You are the project leader for a development team considering developing a software product for tracking blood donations. Market analysis suggests your company could make profits of $125,000 for the first year and then profits decrease by 6% annually. You will sell this software product for 4 years. To cover the software product development costs right now, your company would withdraw money from an investment account paying 7.5% APR compounded annually. Using P/A-G equations, What is the maximum amount of money you will recommend that your company be willing to invest now to develop this software product?
Answer #1

Maximum amount that should be invested = Present value of the funds

= PV of annuity with growth = (P/ (r-g)) * (1- ((1+g)/(1+r))^n)

= (125000/(7.5%+6%))* (1- ((1-0.06)/(1+0.075))^4)=

384606.26